OpenStudy (anonymous):

plz help! Mrs. Brenton has 8 yards of ribbon for her class to make a craft. Each student making the craft will need 1/2 of a yard of ribbon. How many students will be able to make the craft with the amount of ribbon Mrs. Brenton has? A: 8 B: 4 C: 16 D: 2

OpenStudy (anonymous):

8 divided by 1/2 8 * 2/1 = 16 check... 16 * 1/2 = 16/2 = 8 (correct) answer C. 16